# Object-Centric Image to Video Generation with Language Guidance

### v1

@@ 3.3 Video Rendering
$\text{TextOCVP}_{\text{SAVi}}$ Decoder   This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al., 2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$. The object masks are normalized across the slot dimension, and the representations are combined via a weighted sum to render video frames:
@@
Figure 3: Text-conditioned object-centric predictor.
$\text{TextOCVP}_{\text{DINO}}$ Decoder   This decoder variant decodes the object slots in two distinct stages. First, following DINOSAUR (Seitzer et al., 2023), an MLP-based Spatial Broadcast Decoder (Watters et al., 2019) is used to generate object features along with their corresponding masks. Similar to the $\text{TextOCVP}_{\text{SAVi}}$ decoder, the object masks are normalized and combined with the object features in order to reconstruct the encoded features $\hat{\textbf{h}}_{t}\in{\mathbb{R}}^{L\times D_{h}}$. In the second stage, the reconstructed features $\hat{\textbf{h}}_{t}$ are arranged into a grid format and processed with a CNN decoder to generate the corresponding video frame $\hat{\textbf{X}}_{t}$.
@@ 3.4 Training and Inference
Object-Centric Learning   We first train the scene parsing and video rendering modules for decomposing video frames into object-centric representations by minimizing a reconstruction loss. In the $\text{TextOCVP}_{\text{SAVi}}$ variant, these modules are trained simply by reconstructing the input images, whereas in $\text{TextOCVP}_{\text{DINO}}$ they are trained by jointly minimizing an image and a feature reconstruction loss:
$\mathcal{L}_{\text{SAVi}}=\frac{1}{T}\sum_{t=1}^{T}||\hat{\textbf{X}}_{t}-\textbf{X}_{t}||,$ (4)
$\mathcal{L}_{\text{DINO}}=\frac{1}{T}\sum_{t=1}^{T}||\hat{\textbf{X}}_{t}-\textbf{X}_{t}||+||\hat{\textbf{h}}_{t}-\textbf{h}_{t}||.$ (5)
Predictor Training   Given the pretrained scene parsing and rendering modules, we train our TextOCVP predictor for text-conditioned video generation using a dataset containing paired videos and text descriptions. Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual description $\mathcal{C}$, our TextOCVP predictor forecasts the subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$. This process is repeated autoregressively, i.e. the predicted slots are appended to the input in the next time step, in order to generate the set of slots for the subsequent $T$ time steps. This autoregressive training, in contrast to teacher forcing, enforces our predictor to operate with imperfect inputs, leading to better modeling of long-term dynamics at inference time.
Our predictor is trained by minimizing the following combined loss:
$\mathcal{L}_{\text{TextOCVP}}=\frac{1}{T}\sum_{t=2}^{T+1}\lambda_{\text{Img}}\mathcal{L}_{\text{Img}}+\lambda_{\text{Slot}}\mathcal{L}_{\text{Slot}},$ (6)
with $\mathcal{L}_{\text{Img}}=||\hat{\textbf{X}}_{t}-\textbf{X}_{t}||_{2}^{2}\;\;\text{and}\;\;\mathcal{L}_{\text{Slot}}=||\hat{\textbf{S}}_{t}-\textbf{S}_{t}||_{2}^{2},$ (7)
@@
Inference At inference time, TextOCVP receives as input a single reference frame and a language instruction. Our model parses the seed frame into object slots and autoregressively predicts future object states and video frames conditioned on the given textual description. By modifying the language instruction, TextOCVP can generate a new sequence continuation that performs the specified task while preserving a consistent scene composition.
@@ 4.1 Experimental Setup
4.1.1 Datasets
We evaluate the performance of TextOCVP for text-driven image-to-video generation on two different datasets, namely CATER and CLIPort. Further dataset details are provided in Appendix D.
CATER   CATER (Girdhar & Ramanan, 2020) is a dataset that consists of long video sequences, featuring two simultaneously moving 3D objects, with the motion described by a textual caption. We used the CATER-Hard variant introduced by Hu et al. (2022). This dataset contains 30,000 sequences featuring between three and eight objects, two of which follow a predefined action pattern. In our experiments, we resize the images to $64\times 64$.
CLIPort   CLIPort (Shridhar et al., 2022) is a robot manipulation dataset consisting of video-caption pairs. We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task. Each sequence features a table with multiple bowls and blocks of different color, placed at random positions on the table. The corresponding caption describes the action of the robot arm picking a block of a specific color and putting it into a specific bowl.
4.1.2 Baselines
@@
4.1.3 Implementation Details
All our models are implemented in PyTorch and trained on a single NVIDIA A6000 (48Gb) GPU. $\text{TextOCVP}_{\text{SAVi}}$ closely follows Kipf et al. (2022) for the design of the scene parsing and video rendering modules. $\text{TextOCVP}_{\text{DINO}}$ leverages DINOv2 (Oquab et al., 2023) as image encoder, uses a four-layer MLP-based Spatial-Broadcast Decoder (Watters et al., 2019), which is shared among all slots, to decode the object slots into object features and masks. Additionally, it uses a CNN decoder to map the reconstructed scene features back to images. On the CATER dataset, we use the $\text{TextOCVP}_{\text{SAVi}}$ variant with eight 128-dimensional object slots; whereas on CLIPort we employ the $\text{TextOCVP}_{\text{DINO}}$ variant with ten 128-dimensional slots. Our predictor module is an eight-layer transformer with 512-dimensional tokens, eight attention heads, and a hidden dimension of 1024. Further implementation details are provided in Appendix B.
4.2 Results
4.2.1 CATER Results
Table 1: Quantitative evaluation on CATER for prediction horizons of $T=9$ and $T=19$. TextOCVP outperforms all baselines. Best two results are highlighted in boldface and underlined, respectively.
CATER1→9 CATER1→19
[…]
OCVP 0.874 0.078 0.854 0.101
Non-OC TF 0.874 0.092 0.849 0.112
SEER 0.723 0.245 0.535 0.299
MAGE 0.877 0.108 0.871 0.111
TextOCVP 0.924 0.035 0.904 0.042
On the CATER dataset, we train the models to generate nine future frames given a single reference frame and a text caption. In Tab. 1 we report quantitative evaluations on CATER using the same setting as in training, i.e. predicting $T=9$ frames, as well as when predicting $T=19$ future frames. In both settings, TextOCVP outperforms all other models, demonstrating superior perceptual quality.
[…]
@@ 4.2.1 CATER Results
Figure 4: Qualitative evaluation for text-driven image-to-video generation on CATER. Top row shows ground truth frames. TextOCVP generates sharp future frames, whereas MAGE blurs and misses objects.
Fig. 4 shows a qualitative comparison between TextOCVP and MAGE. Our proposed method generates a sequence that is closely aligned to the ground-truth, whereas MAGE predictions feature multiple errors and artifacts, including missing objects, blurry contours, and significant changes on object shapes. This highlights the effectiveness of object-centric representations to accurately represent and model object dynamics. Additional qualitative results are provided in Appendix E.
4.2.2 CLIPort Results
@@
Non-OC TF 0.901 0.184 0.872 0.210
SEER 0.887 0.141 0.622 0.331
$\text{MAGE}_{\text{DINO}}$ 0.940 0.064 0.931 0.075
TextOCVP 0.950 0.062 0.931 0.078
‘put the gray block in the brown bowl
[…]
‘put the blue block in the gray bowl
[…]
Figure 5: Qualitative evaluation on CLIPort. Top rows depict the ground truth frames. TextOCVP successfully completes the pick-and-place task, whereas MAGE fails to predict robot motion (a) or the block disappears (b).
[…]
However, we identify that TextOCVP often generates artifacts in the background and lacks textured details in certain objects, thus degrading its quantitative performance despite generating accurate future frames. Additional qualitative results are provided in Appendix E.
4.3 Model Analysis
4.3.1 Ablation Studies
@@
8 25.95 0.890 0.079
10 26.99 0.950 0.062
We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s image-to-video generation performance. The results are presented in Tab. 6(d).
Number of Layers (Tab. 6(a))   TextOCVP performs best with $N_{\text{Pred}}=8$ predictor layers. Scaling beyond eight layers was not explored, as the performance improvements when increasing from four to eight layers were marginal.
Residual Connection (Tab. 6(c))   Applying a residual connection to the predictor output, $\hat{\textbf{S}}_{t+1}:=\hat{\textbf{S}}_{t+1}+\textbf{S}_{t}$, enhances the model’s performance by improving the temporal consistency of the predicted slots.
Text Encoder (Tab. 6(b))   We evaluate the performance of different text encoders, including a vanilla transformer, a frozen T5 encoder, and a fine-tuned (FT) T5. A frozen T5 module led to the best TextOCVP performance on both CATER and CLIPort datasets. We argue that given the relatively small vocabulary size on CLIPort, fine-tuning T5 did not prove to be beneficial.
Number of Slots (Tab. 6(d))   Using $N_{\mathbf{S}}=10$ slots resulted in better performance for TextOCVP on CLIPort, even though each scene can be described with eight slots (six objects, robot arm and background). This observation suggests that additional slots can be used as registers in order to help with internal computations (Darcet et al., 2024).
Table 4: Quantitative evaluation on a CLIPort test set with unseen colors. We report the video generation performance and its drop relative to the known-color dataset.
@@
4.3.2 Model Robustness
We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set featuring color variations in the text instructions that were not encountered during training. This evaluation measures how well the models generalize to unseen scene features, highlighting their ability to handle novel configurations.
Tab. 4 presents the results on this evaluation set. We report both the video generation performance as well as the performance drop relative to the evaluation set with known colors. Our model outperforms $\text{MAGE}_{\text{DINO}}$ for both prediction horizons. Most notably, TextOCVP demonstrates significantly higher robustness to objects with previously unseen colors compared to the baseline, showing a much smaller drop in performance for the perceptual LPIPS metric. This highlights the effectiveness of leveraging object-centric representations for image-to-video generation and planning, in contrast to relying on holistic scene representations that struggle with novel scene compositions.
4.3.3 Controllability
@@
[…]
Figure 7: Qualitative evaluation of TextOCVP controllability on CATER. Top row shows the ground truth sequence. We underline the changed actions with respect to the original caption. We demonstrate TextOCVP’s controllability by generating multiple sequences from the same reference frame, each conditioned on a different text instruction.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the cyan block in the brown bowl
Original
Caption
Refer to captionRefer to captionRefer to caption‘put the blue block in the brown bowl
Changed
Block
[…]
Figure 8: Qualitative evaluation of TextOCVP controllability on CLIPort. Top row depicts ground-truth frames. TextOCVP correctly generates a sequence where the robot picks up the correct block and places it into the bowl as specified in the textual instruction.
@@
A key objective of text-driven image-to-video generation is to provide control over the model generations. This is achieved through language instructions that describe the objects in the scene and their expected motion. We qualitatively assess the controllability of our model on both the CATER and CLIPort datasets.
@@
CATER   In Fig. 7, we showcase the control that TextOCVP provides over its generations. We qualitatively evaluate this ability by generating multiple videos conditioned on the same reference frame while varying the language instruction. We experiment with altering both the moving and target objects and their actions in the language instruction, as well as providing instructions that include a number of tasks beyond what the model was trained on.
As shown in Fig. 7, TextOCVP successfully identifies the objects described in the text and executes the instructions accordingly. Notably, our model can distinguish between two nearly identical purple cones in the scene, despite their identical shapes and color, and correctly generates a sequence with the specified motion.
@@
This demonstrates that key object attributes, such as size or color, are effectively captured through the text-to-slot attention mechanism. Leveraging this information, the text-conditioned predictor accurately predicts the future motion of each individual object as specified in the instruction. These results highlight the benefits of combining object-centric representations with text-based conditioning.
CLIPort   We perform a similar experiment on the CLIPort dataset, as shown in Fig. 8. A single frame containing multiple colored blocks and bowls is provided to TextOCVP. By modifying the textual instruction, we can control which block the robot arm picks up and the bowl where it is placed. In both examples, TextOCVP successfully selects the correct block and places it in the specified bowl. Furthermore, the movement of the robot arm adapts to the motion described in the instruction, demonstrating accurate and responsive control.
@@ 5 Conclusion
We presented TextOCVP, a novel object-centric model for text-driven image-to-video generation. Given a single input image and a text description, TextOCVP generates a sequence that matches the description by parsing the environment into object representations and modeling their dynamics conditioned on the textual instruction. This is achieved with a novel text-conditioned object-centric transformer, which predicts future object states by modeling the spatio-temporal relationships between objects while incorporating the guidance from text. Through extensive evaluations, we demonstrated that TextOCVP outperforms existing text-conditioned models for the image-to-video generation, highlighting our model’s ability to generate consistent long sequences and adapt its predictions based on language input. Moreover, we validated our architectural choices through ablation studies, highlighting the importance of combining textual and object-centric information. With its structured latent space and superior controllability, TextOCVP offers potential for robotics applications, where it can serve as a controllable world model, enabling efficient planning, reasoning, and decision-making.

### v2

@@ 3.3 Video Rendering
$\text{TextOCVP}_{\text{SAVi}}$ Decoder
This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al. 2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$. The object masks are normalized across the slot dimension, and the representations are combined via a weighted sum to render video frames:
@@
Figure 3: Overview of our text-conditioned object-centric video predictor. Our predictor employs a transformer-based architecture to model the spatio-temporal dynamics of object slots via self-attention, while incorporating textual guidance through a text-to-slot cross-attention mechanism.
$\text{TextOCVP}_{\text{DINO}}$ Decoder
This decoder variant decodes the object slots in two distinct stages. First, following DINOSAUR (Seitzer et al. 2023), an MLP-based Spatial Broadcast Decoder (Watters et al. 2019) is used to generate object features along with their corresponding masks. Similar to the $\text{TextOCVP}_{\text{SAVi}}$ decoder, the object masks are normalized and combined with the object features in order to reconstruct the encoded features $\hat{\textbf{h}}_{t}\in{\mathbb{R}}^{L\times D_{h}}$. In the second stage, the reconstructed features $\hat{\textbf{h}}_{t}$ are arranged into a grid format and processed with a CNN decoder to generate the corresponding video frame $\hat{\textbf{X}}_{t}$.
@@ 3.4 Training and Inference
Object-Centric Learning
We first train the scene parsing and video rendering modules for parsing video frames into object-centric representations by minimizing a reconstruction loss. In the $\text{TextOCVP}_{\text{SAVi}}$ variant ($\mathcal{L}_{\text{SAVi}}$), these modules are trained by reconstructing the input images, whereas in $\text{TextOCVP}_{\text{DINO}}$ ($\mathcal{L}_{\text{DINO}}$) they are trained by jointly minimizing an image and a feature reconstruction loss:
[…]
Predictor Training
Given the pretrained scene parsing and rendering modules, we train our TextOCVP predictor for text-conditioned video prediction using a dataset containing paired videos and text descriptions. Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual embeddings $\mathbf{C}$, our TextOCVP predictor forecasts the subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$. This process is repeated autoregressively, i.e. the predicted slots are appended to the input in the next time step, in order to generate the set of slots for the subsequent $T$ time steps. This autoregressive training, in contrast to teacher forcing, enforces our predictor to operate with imperfect inputs, leading to better modeling of long-term dynamics at inference time.
Our predictor is trained by minimizing the following loss:
$\mathcal{L}_{\text{TextOCVP}}=\frac{1}{T}\sum_{t=2}^{T+1}\left(\lambda_{\text{Img}}\mathcal{L}_{\text{Img}}+\lambda_{\text{Slot}}\mathcal{L}_{\text{Slot}}\right),$ (6)
[…]
@@
Inference
At inference time, TextOCVP receives as input a single reference frame and a language instruction. Our model parses the seed frame into object slots and autoregressively predicts future object states and video frames conditioned on the given textual description. By modifying the language instruction, TextOCVP can generate a new sequence continuation that performs the specified task while preserving a consistent scene composition.
@@ 4.1 Experimental Setup
Datasets
We evaluate TextOCVP for text-conditioned video prediction on the CATER and CLIPort datasets:
CATER (Girdhar and Ramanan 2020) is a dataset of long video sequences showing 3D objects in motion, each accompanied by a descriptive caption. We used the CATER-hard variant (Hu, Luo, and Chen 2022), containing 30,000 sequences with $64\times 64$ frames featuring between two and eight objects, two of which follow the scripted motion.
CLIPort (Shridhar, Manuelli, and Fox 2022) is a robot manipulation dataset featuring video-caption pairs. We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task, where the caption describes placing a block into a bowl.
Baselines
@@
Refer to captiont=1t=122101020203030MAGE TextOCVP
@@
Figure 4: Text-guided video prediction result on CATER. Top row shows ground truth. TextOCVP predicts sharp and accurate frames, whereas MAGE blurs and misses objects.
Implementation Details
All models are implemented in PyTorch and trained on a single NVIDIA A6000 (48Gb) GPU. $\text{TextOCVP}_{\text{SAVi}}$ closely follows Kipf et al. (2022) for scene parsing and video rendering. $\text{TextOCVP}_{\text{DINO}}$ uses DINOv2 (Oquab et al. 2024) as image encoder, a four-layer MLP Spatial-Broadcast Decoder (Watters et al. 2019) to decode the slots into object features and masks, and a CNN decoder to map the reconstructed scene features back to images. On CATER, we use the $\text{TextOCVP}_{\text{SAVi}}$ variant with eight 128-dimensional object slots, whereas on CLIPort we employ the $\text{TextOCVP}_{\text{DINO}}$ variant with ten 128-dimensional slots. Our predictor module is an eight-layer transformer with 512-dimensional tokens, eight attention heads, and a hidden dimension of 1024.
Further experiment details are provided in Appendix B-E.
4.2 Results
CATER Results
On CATER, we train the models to predict nine future frames given a single reference frame and a text caption. In Table 1(a) we report quantitative evaluations on CATER using the same setting as in training, i.e. predicting $T=9$ frames, as well as when predicting $T=19$ future frames. In both settings, TextOCVP outperforms all other models, demonstrating superior perceptual quality.
Fig. 4 qualitatively compares TextOCVP and MAGE. Our proposed method generates a sequence that is closely aligned to the ground-truth, whereas MAGE predictions feature multiple errors and artifacts, including missing objects, blurry contours, and significant changes on object shapes. This highlights the effectiveness of object-centric representations to accurately represent and model object dynamics.
@@ 4.2 Results
(a) CATER Evaluation
CATER1→9 CATER1→19
[…]
OCVP 0.874 0.078 4.16 0.854 0.101 8.08
Non-OC 0.874 0.092 3.04 0.849 0.112 8.62
SEER 0.723 0.245 11.23 0.535 0.299 17.29
MAGE 0.877 0.108 3.46 0.871 0.111 5.88
TextOCVP 0.922 0.036 2.16 0.902 0.044 5.09
(b) CLIPort Evaluation
@@
Non-OC 0.901 0.184 8.13 0.872 0.210 13.23
SEER 0.887 0.141 6.80 0.622 0.331 8.29
$\text{MAGE}_{\text{DINO}}$ 0.940 0.064 2.11 0.931 0.075 2.59
TextOCVP 0.950 0.062 1.36 0.931 0.078 2.23
@@
8 0.946 0.079 5.62
10 0.950 0.062 1.36
CLIPort Results
‘put the blue block in the gray bowl
[…]
TextOCVPt=1t=12020303040405050Missing
[…]
Nevertheless, we identify that TextOCVP often generates visual artifacts and lacks fine-grained textures, which can affect its quantitative scores despite accurate predictions.
Further figures and results are provided in Appendix F.
4.3 Model Analysis
Ablation Studies
We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s video prediction performance. The results are presented in Table 2(d).
Number of Layers (Table 2(a))   TextOCVP performs best with $N_{\text{P}}=8$ predictor layers, outperforming shallower variants. We did not explore scaling beyond eight layers due to the significant increase in the number of parameters.
Residual Connection (Table 2(c))   Incorporating a residual connection in the predictor, $\hat{\textbf{S}}_{t+1}:=\hat{\textbf{S}}_{t+1}+\textbf{S}_{t}$, enhances the model’s performance by improving the temporal consistency of the predicted slots.
Text Encoder (Table 2(b))   We evaluate the performance of different text encoders, including a vanilla transformer, a frozen T5 encoder, and a fine-tuned (FT) T5. A frozen T5 module led to the best TextOCVP performance on the CLIPort dataset. Although all encoders led to similar performance on the CATER dataset, we opted for the frozen T5 to avoid unnecessarily increasing the model’s parameter count.
Number of Slots (Table 2(d))   Using $N_{\mathbf{S}}=10$ slots resulted in better performance for TextOCVP on CLIPort, even though each scene can be described with eight slots (six objects, robot arm and background). This observation suggests that additional slots can be used as registers in order to help with internal computations (Darcet et al. 2024).
Table 3: Quantitative evaluation on a CLIPort test set with unseen colors. We report the video prediction performance and its drop relative to the known-color dataset.
@@ 4.3 Model Analysis
Model Robustness
@@
t=1t=12215153030
Figure 6: Evaluation of TextOCVP’s controllability on CATER. Top row shows the ground truth sequence. TextOCVP enables predicting different sequence continuations conditioned on distinct text instructions.
We evaluate the generalization ability of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set featuring color variations in the text instructions that were not encountered during training. This setup assesses the models’ ability to generalize to novel scene configurations and adapt to unseen text-visual combinations.
Table 3 presents the results on this evaluation set with unseen colors. We report both the video prediction performance as well as the performance drop relative to the evaluation set with known colors. Our model consistently outperforms $\text{MAGE}_{\text{DINO}}$ for both prediction horizons. Most notably, TextOCVP demonstrates significantly higher robustness to objects with previously unseen colors compared to the baseline, showing a much smaller drop in performance for the perceptual LPIPS metric. These results highlight the advantages of structured object-centric representations, which provide greater generalization and robustness in video prediction—particularly in contrast to holistic scene representations that struggle with novel scene compositions.
Controllability
@@
A key objective of text-driven video prediction is to provide fine-grained control over the prediction process, which is enabled through language instructions describing the objects in the scene and their expected motion.
@@
In Fig. 6, we qualitatively showcase the level of control that TextOCVP provides over its generations on the CATER dataset. We qualitatively evaluate this ability by generating multiple video continuations from the same reference frame while varying the language instruction. These variations include changes to both the target objects and their actions, as well as providing instructions that include a number of actions beyond what the model was trained on. As shown in Fig. 6, TextOCVP successfully identifies the relevant objects and executes the actions described in the text instructions. Notably, our model distinguishes between two nearly identical purple cones, despite the identical shapes and color, and correctly generates a sequence with the specified motion.
@@
These results show that key object attributes, such as size or color, are effectively captured through the text-to-slot attention mechanism, enabling the predictor to accurately forecast the motion of each individual object. This highlights the benefits of combining object-centric representations with text guidance for fine-grained, controllable video prediction.
Interpretability
Refer to captiont=1t=1Refer to caption55Refer to caption1010Refer to caption2020Refer to caption3030
‘the cone is sliding to (-3, -3).
the snitch is picked up and placed to (-1, 2 ).’
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 3
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 4
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 6
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Figure 7: Text-to-slot attention weights for different objects in a video. Slots representing objects described in the caption focus on relevant text tokens, such as target locations.
In Fig. 7, we visualize the text-to-slot attention weights—averaged across attention heads—for different slots in a CATER sequence, revealing how textual information guides model predictions. We observe that slots corresponding to objects described in the textual instruction focus on relevant text tokens, such as target coordinate locations. These results demonstrate that the text-to-slot attention mechanism effectively aligns the textual information with the object-centric representations, enabling the model to generate accurate video sequences based on the given text instructions, while providing superior interpretability.
@@ 5 Conclusion
We presented TextOCVP, a novel object-centric model for text-conditioned video prediction. Given a single input image and a text description, TextOCVP generates frames by parsing the scene into object representations and modeling their dynamics conditioned on the textual instruction. This is accomplished through a text-conditioned object-centric transformer, which predicts future object states by modeling the spatio-temporal relationships between objects while incorporating textual guidance. Through extensive evaluations, we demonstrated that TextOCVP outperforms existing models for text-driven video prediction from a single frame, highlighting our model’s ability to predict over long prediction horizons and adapt its predictions based on the provided textual description. Moreover, we validated our architectural choices through ablation studies, highlighting the importance of combining textual and object-centric information, as well as demonstrating TextOCVP’s robustness and interpretability. With its structured latent space and superior controllability, TextOCVP offers potential for applications in robotics, where it can serve as a controllable world model, enabling efficient planning, reasoning, and decision-making.